# Nuclear chiral density wave in neutron stars?

### v1

@@ 1 Introduction: Chiral density wave in dense matter
The study of neutron star properties is closely tied to the physics of dense matter governed by Quantum Chromodynamics (QCD). It is an open question whether the densities in neutron stars are large enough for quark matter to exist in its core, although some indications for a deconfinement transition have been pointed out [1]. A transition between nuclear and quark matter inside the star would provide a natural environment for a spatially inhomogeneous mixed phase [2], not unlike the nuclear pasta phases conjectured in the crust-core transition of the star [3]. Besides this conceptually straightforward phase separation, more subtle effects that break rotational and/or translational invariance on a microscopic level are possible [4, 5]. Thermal fluctuations tend to disfavor such phases, but isolated neutron stars have essentially zero temperature compared to the relevant baryon chemical potentials. Similarly, one may ask whether neutron star conditions such as electric charge neutrality and equilibrium with respect to the weak interactions tend to favor or disfavor anisotropic phases, see Fig. 1. We shall address this question for the chiral density wave (CDW), a certain anisotropic phase that is particularly simple to implement theoretically and that is conjectured to exist in the vicinity of the chiral phase transition [6].
@@
The CDW is induced by a chiral condensate that rotates periodically between the scalar and pseudoscalar sectors along a certain spatial direction. Unless rotational symmetry is already broken explicitly, for example by a magnetic field, the CDW breaks rotational symmetry spontaneously. The CDW has been studied before within phenomenological models, mainly based on quark degrees of freedom, like the Nambu–Jona-Lasinio model [7] or the quark-meson model [8], and first applications to neutron stars were considered [9]. Much less work has been devoted to the CDW in nuclear matter, where a few studies exist [10, 11], however without systematic explorations of astrophysical relevance. Moreover, many of these studies were restricted either to simple approximations that completely ignored nucleonic vacuum fluctuations or plagued by artifacts induced by these fluctuations, which seemed to arise in conjunction with the CDW. Astrophysical applications and the theoretical issues were both addressed recently in a nucleon-meson model that has a well-defined chiral limit [12, 13]. In these proceedings we review the main results of these two works.
@@ 2 Model: Free energy and parameter fit
The nucleon-meson model we work with contains neutrons and protons as well as scalar and vector mesons and Yukawa interactions between the nucleons and mesons. Details can be found for instance in Refs. [12, 13, 14, 15], with extensions to include strangeness discussed in Ref. [16]. The CDW ansatz is implemented in the expectation values of the sigma and the neutral pion,
@@
with the pion decay constant $f_{\pi}$, parameters $a_{1},\ldots,a_{4}$ and $\epsilon$ (the latter giving rise to a nonzero pion mass), and the vacuum mass of the nucleon $m_{N}$. We have separated the part that is only nonzero in the presence of a CDW,
@@
and where $\ell$ is a renormalization scale, which we choose to have the form [12]
@@
For our results we will choose the numerical constant $c=1/\exp(1+2\pi^{2}/g_{\sigma}^{2})\simeq 0.3$, which is the limit below which the effective potential becomes unbounded with respect to $q$. Curing this unboundedness is the main reason for the choice (9), which also cures unphysical phenomena of the CDW at large densities for $c=1$ [12]. Choosing the smallest value for $c$ renders the CDW as energetically favorable as possible, but re-introduces an unphysical boundary of the anisotropic phase at high densities. This choice makes our main result – the absence of the CDW in neutron stars – robust in the sense that no other choice of $c$ that respects boundedness of the potential in the $q$ direction will change our conclusion. Finally, the vector meson potential in Eq. (4) is
@@
The model parameters are chosen to reproduce vacuum properties and properties of isospin-symmetric nuclear matter at saturation, see Ref. [13] for a detailed explanation of the fitting procedure. Two of the quantities used for the fit, the (Dirac) mass of the nucleon at saturation $M_{0}$ and the incompressibility $K$, have significant empirical uncertainties. In the main results we shall choose $K=300\,{\rm MeV}$, which is at the upper boundary (perhaps somewhat beyond) of the empirical range, following a similar logic as explained for the choice of the constant $c$: as will become clear from the results, our main qualitative conclusions are robust against making nuclear matter softer by choosing a smaller value for $K$. The mass $M_{0}$ will be varied to explore the parameter space of the model. (When $M_{0}$ is varied, the parameters of the model are adjusted such that the physical properties of the vacuum and at saturation are maintained.)
Importantly, we also employ pure neutron matter to fix our parameters [17]. This concerns mostly the choice of the vector meson self-couplings. We shall work with two different fits,
@@
Their different effect on pure neutron matter is shown in Fig. 2, compared with results from chiral effective theory [18]. Using the single degree of freedom $d$ from each of the fits (11) to reproduce the binding energy $E_{0}|_{0.5n_{0}}\simeq 10$ MeV, we see that only Fit(00d) is able to match the binding energy over a large density range, excellent agreement with effective field theory being reached for $M_{0}=0.75m_{N}$.
Figure 2: Binding energy of pure neutron matter as a function of baryon density $n_{B}$ normalized to saturation density $n_{0}$, compared with results from chiral effective field theory (green band).
@@ 3 Results: CDW in neutron stars?
Let us first study the effect of the neutron star conditions on the fate of the CDW. This is done in Fig. 3. In the chiral limit of vanishing pion mass $m_{\pi}=0$, we see that the CDW phase may not exist at all (towards small $M_{0}$) or appears in an intermediate region according to the schematic expectation of Fig. 1 (towards large $M_{0}$), with this intermediate region increasing as $M_{0}$ is increased. This trend is maintained in the case of a physical pion mass, where the phase transition without CDW is a crossover as a consequence of including the Dirac sea [12], and where the transition from isotropic nuclear matter to the CDW is of first order. There is a significant difference between the two fits (11), with Fit(ddd) being more favorable for the CDW. Most importantly, we see that neutron star conditions – similar to the effect of the pion mass – render the CDW less preferred by shifting its onset to larger chemical potentials. In the left panel (upper right corner) we observe the unphysical boundary of the CDW mentioned above (where, in the present approach no other phase takes over consistently). A different renormalization scale, using for example $c=1$, would fix this, but make the region where the CDW is preferred even smaller.
@@
In the upper right panel, the mass of the most massive star for both fits is compared with observations of PSR J0740+6620 [19] and PSR J0952+0607 [20], two of the heaviest observed compact stars. The only stable stars with a CDW core are given by the segment on the (blue) Fit(00d) curve to the right of the dot. They are far from meeting the astrophysical constraints. For Fit(ddd), we see that for all $M_{0}$ considered here the equation of state is too soft to create sufficiently heavy stars (despite using an incompressibility at the upper end of the empirical range). In contrast, Fit(00d) does produce realistically heavy stars, which validates the model, albeit in a regime that does not exhibit a CDW. (For a specific set of model parameters needed to create two-solar-mass stars, for example with $M_{0}=0.76\,m_{N}$, see Ref. [13].) These realistic stars are hybrid in the sense that they contain a crossover to chirally restored matter. This is suggested by the pale dashed lines in the upper left panel, which indicate the chiral transition in the chiral limit. The crossover has a particularly striking effect in the $\mu_{n}$ curve of Fit(ddd), where the discontinuity results from a twin star configuration, i.e., a mass-radius curve that shows stars with two different radii for the same mass (a feature usually assumed to be only possible in the presence of first-order transitions).
@@ 4 Summary and outlook
An obvious question for future work is whether a magnetic field can stabilize the CDW, as suggested by previous work in other models [21, 22], and if yes whether the necessary magnetic field strengths are realistic for neutron star interiors. One should also check the model dependence of our results, for instance by employing the extended linear sigma model of Refs. [11, 23, 24]. Our results depend strongly on whether the nucleonic vacuum fluctuations are taken into account, and neither completely dropping them (previous work [11]) as well as including them without further improvements of the approach (work reviewed here) are controlled approximations due to the large values of the Yukawa couplings. A useful complementary approach would therefore be to include the CDW in a controlled strong-coupling calculation based on the gauge-gravity duality, where it is known how to describe nuclear matter (with and without isospin-asymmetry) [25, 26].

### v2

@@ 1 Introduction: Chiral density wave in dense matter
The study of neutron star properties is closely tied to the physics of dense matter governed by Quantum Chromodynamics (QCD). It is an open question whether the densities in neutron stars are large enough for quark matter to exist in its core, although some indications for a deconfinement transition have been pointed out [Annala:2019puf]. A transition between nuclear and quark matter inside the star would provide a natural environment for a spatially inhomogeneous mixed phase [Schmitt:2020tac], not unlike the nuclear pasta phases conjectured in the crust-core transition of the star [Ravenhall:1983uh]. Besides this conceptually straightforward phase separation, more subtle effects that break rotational and/or translational invariance on a microscopic level are possible [Alford:2000ze, Buballa:2014tba]. Thermal fluctuations tend to disfavor such phases, but isolated neutron stars have essentially zero temperature compared to the relevant baryon chemical potentials. Similarly, one may ask whether neutron star conditions such as electric charge neutrality and equilibrium with respect to the weak interactions tend to favor or disfavor anisotropic phases, see Fig. 1. We shall address this question for the chiral density wave (CDW), a certain anisotropic phase that is particularly simple to implement theoretically and that is conjectured to exist in the vicinity of the chiral phase transition [Nickel:2009wj].
@@
The CDW is induced by a chiral condensate that rotates periodically between the scalar and pseudoscalar sectors along a certain spatial direction. Unless rotational symmetry is already broken explicitly, for example by a magnetic field, the CDW breaks rotational symmetry spontaneously. The CDW has been studied before within phenomenological models, mainly based on quark degrees of freedom, like the Nambu–Jona-Lasinio model [Nakano:2004cd] or the quark-meson model [Carignano:2014jla], and first applications to neutron stars were considered [Buballa:2015awa]. Much less work has been devoted to the CDW in nuclear matter, where a few studies exist [Dautry:1979bk, Heinz:2013hza], however without systematic explorations of astrophysical relevance. Moreover, many of these studies were restricted either to simple approximations that completely ignored nucleonic vacuum fluctuations or plagued by artifacts induced by these fluctuations, which seemed to arise in conjunction with the CDW. Astrophysical applications and the theoretical issues were both addressed recently in a nucleon-meson model that has a well-defined chiral limit [Pitsinigkos:2023xee, Papadopoulos:2024agt]. In these proceedings we review the main results of these two works.
@@ 2 Model: Free energy and parameter fit
The nucleon-meson model we work with contains neutrons and protons as well as scalar and vector mesons and Yukawa interactions between the nucleons and mesons. Details can be found for instance in Refs. [Pitsinigkos:2023xee, Papadopoulos:2024agt, Drews:2014spa, Fraga:2018cvr], with extensions to include strangeness discussed in Ref. [Fraga:2022yls]. The CDW ansatz is implemented in the expectation values of the sigma and the neutral pion,
@@
with the pion decay constant $f_{\pi}$, parameters $a_{1},\ldots,a_{4}$ and $\epsilon$ (the latter giving rise to a nonzero pion mass), and the vacuum mass of the nucleon $m_{N}$. All parameters are renormalized at this point. For the details of the renormalization procedure (with the help of proper time regularization), starting from the underlying Lagrangian with bare parameters, see Ref. [Pitsinigkos:2023xee]. We have separated the part that is only nonzero in the presence of a CDW,
@@
and where $\ell$ is a renormalization scale, which we choose to have the form [Pitsinigkos:2023xee]
@@
For our results we will choose the numerical constant $c=1/\exp(1+2\pi^{2}/g_{\sigma}^{2})\simeq 0.3$, which is the limit below which the effective potential becomes unbounded with respect to $q$. Curing this unboundedness is the main reason for the choice (9), which also cures unphysical phenomena of the CDW at large densities for $c=1$ [Pitsinigkos:2023xee]. Choosing the smallest value for $c$ renders the CDW as energetically favorable as possible, but re-introduces an unphysical boundary of the anisotropic phase at high densities. This choice makes our main result – the absence of the CDW in neutron stars – robust in the sense that no other choice of $c$ that respects boundedness of the potential in the $q$ direction will change our conclusion. Finally, the vector meson potential in Eq. (4) is
@@
The model parameters are chosen to reproduce vacuum properties and properties of isospin-symmetric nuclear matter at saturation, see Ref. [Papadopoulos:2024agt] for a detailed explanation of the fitting procedure. Two of the quantities used for the fit, the (Dirac) mass of the nucleon at saturation $M_{0}$ and the incompressibility $K$, have significant empirical uncertainties. In the main results we shall choose $K=300\,{\rm MeV}$, which is at the upper boundary (perhaps somewhat beyond) of the empirical range, following a similar logic as explained for the choice of the constant $c$: as will become clear from the results, our main qualitative conclusions are robust against making nuclear matter softer by choosing a smaller value for $K$. The mass $M_{0}$ will be varied to explore the parameter space of the model. (When $M_{0}$ is varied, the parameters of the model are adjusted such that the physical properties of the vacuum and at saturation are maintained.)
Importantly, we also employ pure neutron matter to fix our parameters [Alford:2022bpp]. This concerns mostly the choice of the vector meson self-couplings. We shall work with two different fits,
@@
Their different effect on pure neutron matter is shown in Fig. 2, compared with results from chiral effective theory [Tews:2018kmu]. Using the single degree of freedom $d$ from each of the fits (11) to reproduce the binding energy $E_{0}|_{0.5n_{0}}\simeq 10$ MeV, we see that only Fit(00d) is able to match the binding energy over a large density range, excellent agreement with effective field theory being reached for $M_{0}=0.75m_{N}$.
Figure 2: Binding energy of pure neutron matter as a function of baryon density $n_{B}$ normalized to saturation density $n_{0}$, compared with results from chiral effective field theory given by the green error band [Tews:2018kmu].
@@ 3 Results: CDW in neutron stars?
Let us first study the effect of the neutron star conditions on the fate of the CDW. This is done in Fig. 3. In the chiral limit of vanishing pion mass $m_{\pi}=0$, we see that the CDW phase may not exist at all (towards small $M_{0}$) or appears in an intermediate region according to the schematic expectation of Fig. 1 (towards large $M_{0}$), with this intermediate region increasing as $M_{0}$ is increased. This trend is maintained in the case of a physical pion mass, where the phase transition without CDW is a crossover as a consequence of including the Dirac sea [Pitsinigkos:2023xee], and where the transition from isotropic nuclear matter to the CDW is of first order. There is a significant difference between the two fits (11), with Fit(ddd) being more favorable for the CDW. Most importantly, we see that neutron star conditions – similar to the effect of the pion mass – render the CDW less preferred by shifting its onset to larger chemical potentials. In the left panel (upper right corner) we observe the unphysical boundary of the CDW mentioned above (where, in the present approach no other phase takes over consistently). A different renormalization scale, using for example $c=1$, would fix this, but make the region where the CDW is preferred even smaller.
@@
In the upper right panel, the mass of the most massive star for both fits is compared with observations of PSR J0740+6620 [Fonseca:2021wxt] and PSR J0952+0607 [Romani:2022jhd], two of the heaviest observed compact stars. The only stable stars with a CDW core are given by the segment on the (blue) Fit(00d) curve to the right of the dot. They are far from meeting the astrophysical constraints. For Fit(ddd), we see that for all $M_{0}$ considered here the equation of state is too soft to create sufficiently heavy stars (despite using an incompressibility at the upper end of the empirical range). In contrast, Fit(00d) does produce realistically heavy stars, which validates the model, albeit in a regime that does not exhibit a CDW. (For a specific set of model parameters needed to create two-solar-mass stars, for example with $M_{0}=0.76\,m_{N}$, see Ref. [Papadopoulos:2024agt].) These realistic stars are hybrid in the sense that they contain a crossover to chirally restored matter. This is suggested by the pale dashed lines in the upper left panel, which indicate the chiral transition in the chiral limit. The crossover has a particularly striking effect in the $\mu_{n}$ curve of Fit(ddd), where the discontinuity results from a twin star configuration, i.e., a mass-radius curve that shows stars with two different radii for the same mass (a feature usually assumed to be only possible in the presence of first-order transitions).
@@ 4 Summary and outlook
An obvious question for future work is whether a magnetic field can stabilize the CDW, as suggested by previous work in other models [Ferrer:2019zfp, Gyory:2022hnv], and if yes whether the necessary magnetic field strengths are realistic for neutron star interiors. Also, one can include the competing/coexisting effect of Cooper pairing and allow for phases that break translational invariance. One should also check the model dependence of our results, for instance by employing the extended linear sigma model of Refs. [Heinz:2013hza, Haber:2014ula, Takeda:2018ldi]. Our results depend strongly on whether the nucleonic vacuum fluctuations are taken into account, and neither completely dropping them (previous work [Heinz:2013hza]) as well as including them without further improvements of the approach (work reviewed here) are controlled approximations due to the large values of the Yukawa couplings. A useful complementary approach would therefore be to include the CDW in a controlled strong-coupling calculation based on the gauge-gravity duality, where it is known how to describe nuclear matter (with and without isospin-asymmetry) [Kovensky:2021ddl, Ecker:2025sjb].